Two planes leave from the same airport at the same time but travel in opposite directions. One flies 60 mi/h faster than the other. They are 1,820 miles apart after 2 hours. How fast is each plane traveling?
Yes, I'm very new to algebra. Thanx for any help!

Answer by stanbon(48558)

Two planes leave from the same airport at the same time but travel in opposite directions. One flies 60 mi/h faster than the other. They are 1,820 miles apart after 2 hours. How fast is each plane traveling?
------------------------
1st plane DATA:
Rate = x mph ; time = 2 hrs ; distance = 2x miles
-----------------
2nd plane DATA:
Rate = x+60 mph ; time = 2 hrs ; distance = 2x+120 miles
----------------
EQUATION:
distance + distance = 1820 miles
2x + 2x+120 = 1820
4x = 1700
x = 425 mph (1st plane rate)
x+60 = 485 mph (2nd plane rate)
============================
Cheers,
Stan H.